# Inference for concave distribution functions under measurement error

### v1

@@ Abstract
We propose an estimator of a concave cumulative distribution function under the measurement error model, where the non-negative variables of interest are perturbed by additive independent random noise. The estimator is defined as the least concave majorant on the positive half-line of the deconvolution estimator of the distribution function. We show its uniform consistency and its square root convergence in law in $\ell_{\infty}(\mathbb{R})$. To assess the validity of the concavity assumption, we construct a test for the nonparametric null hypothesis that the distribution function is concave on the positive half-line, against the alternative that it is not. We calibrate the test using bootstrap methods. The theoretical justification for calibration led us to establish a bootstrap version of Theorem 1 in Söhl and Trabs (2012), a Donsker-type result from which we obtain, as a special case, the limiting behavior of the deconvolution estimator of the distribution function in a bootstrap setting with measurement error. Combining this Donsker-type theorem with the functional delta method, we show that the test statistic and its bootstrap version have the same limiting distribution under the null hypothesis, whereas under the alternative the bootstrap statistic is stochastically smaller. Consequently, the power of the test tends to one, for any fixed alternative, as the sample size tends to infinity. In addition to the theoretical results for the estimator and the test, we investigate their finite-sample performance in simulation studies.
@@ 1 Introduction
Statistical inference under shape restrictions exploits nonparametric assumptions on the shape of the distribution of interest to improve estimates. Additionally, it deals with nonparametric tests about such shape constraints. The books [25, 14, 29] provide an overview of the issues addressed in the context of statistical inference under shape constraints. The results depend on the methodology and on the type of assumption, see for instance [13, 21, 8, 26, 16, 18, 17]. One of the most commonly used constraint, starting from the seminal work of [13], is the concavity of the cumulative distribution function [5, 3], corresponding to a decreasing density [14, Chapter 2.2].
@@
This paper combines theories on deconvolution estimation (i.e., estimation under measurement error) and on estimation and testing under shape constraints. As a shape-constrained estimator, we consider the least concave majorant of the classic kernel deconvolution estimator of the distribution function $F$ of $X$. We show its uniform consistency under general assumptions, and we establish a Donsker-type result for it, i.e. we show that the difference between the estimator and the true concave distribution function converges in law at the rate $\sqrt{n}$ to a centered Borel random variable in $\ell_{\infty}(\mathbb{R})$. Then, inspired by [10, 8], we define our test statistic as the supremum distance between the two estimators. Although the construction is similar, the proof of the theoretical properties is much more involded in the model with measurement error than in those considered in [10, 8]. In particular, we do not have access to a least favorable hypothesis (see Section 5.2) so the calibration of the test requires challenging efforts. The performance of the constrained estimator and that of the test are illustrated through a simulation study. he simulations suggest that, for sufficiently large sample sizes and a moderate noise-to-signal ratio, (i) when $F$ is concave, the constrained estimator outperforms the unconstrained one; (ii) under concavity of $F$, the rejection rate of the test does not exceed the nominal level; and (iii) the test can detect nonconcavity of $F$ even in settings where the distribution of $Y$ is concave.
@@
    a bootstrap version of the Donsker Theorem 1 in [30] from which, as a particular case, one can derive the limit behavior of the deconvolution estimator of the distribution function in a bootstrap setting, under measurement error, see Theorem 7 and (14) below;
@@
The paper is organized as follows. Section 2 provides the model, definitions and assumptions. The constrained estimator is studied in Section 3, which also contains new results about uniform convergence of least concave majorants of continuously differentiable functions, as well as that of their slopes. The test statistic is defined in Section 4, where its limit behavior is studied under both the null and the alternative hypotheses. A bootstrap calibration of the test is given in Section 5, which also contains an alternative calibration that provides a less powerfull test, compared to the bootstrap, but is easier to implement. The main part of the paper ends with Section 6, which is devoted to simulations. The main sections are followed by appendices that are organized as follows. The results from Sections 2 and 3 are proved in Section A. Theorem 5, that provides the limit behavior of the test statistic, is proved in B. Section C provides a generalization of the Donsker result in [30] to a bootstrap setting. The corresponding proofs are given in Section D. Section E provides all the proofs for the bootstrap calibration.
@@ 2.2 Model, objective and basic estimators
Our aim is to build a test for the null hypothesis
@@
against the alternative hypothesis $\mathcal{H}_{1}:\ ``F\text{ is not concave on }[0,\infty)$”. Moreover, if the null hypothesis is accepted, we wish to construct an estimator of $F$ that takes it into account, i.e. a concave estimator. Defining the concave estimator and the test requires to first define Fourier transforms, which we will now do.
@@ 2.3 Assumptions
A key ingredient in our proofs is Theorem 1 in Söhl and Trabs (2012). This is a Donsker-type result that gives, as a special case, the limiting behavior of the deconvolution estimator of the distribution function in the measurement error model. Assumptions K, D and E below are taken from that theorem.
@@ 3.1 Definition of the concave estimator and consistency
The following theorem, which is proved in Section A, provides a general result on convergence of least concave majorants (and their slopes) of a sequence of continuously differentiable functions whose sequence of slopes uniformly converge.
@@
As a corollary (also proved in Section A) of Theorem 1, we derive uniform convergence of our concave estimators.
@@ 3.2 Limit distribution
Theorem 2.
[30] Assume conditions D and E with $\beta<1/2$, $\alpha+3\gamma>2\beta+1$ for some $\gamma\in(\beta,1/2)$ and K with $L=\left\lfloor\alpha+\gamma\right\rfloor$. Furthermore, let $h_{n}^{2\alpha+2\gamma}n\rightarrow 0$ and $h_{n}^{\rho}n\rightarrow\infty$ for some $\rho>4\beta-4\gamma+2$. Then, we have
@@
We are now in position to state the main result of the section, which gives the limit distribution of the concave estimator $\mathcal{M}F_{n}^{d}$. The theorem shows that the estimator converges to the least concave majorant of $F$ at the rate $\sqrt{n}$, so, in particular, the estimator converges to $F$ if the latter is concave on its support. More precisely, the difference between the estimator and the least concave majorant of $F$ converges at the rate $\sqrt{n}$ to a centered Gaussian Borel variable in $\ell_{\infty}(\mathbb{R})$.
@@
$\sqrt{n}(\mathcal{M}F_{n}^{d}-\mathcal{M}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}(0)\quad in\ \ell_{\infty}(\mathbb{R}).$ (9)
@@ 5.1 Bootstrap calibration
We turn to the construction of the bootstrap observations. Fix $m\ll n$ that tends to infinity as $n\to\infty$, and conditionally on the original sample $Y_{1},\dots,Y_{n}$, consider a sample of $m$ iid random variables $X_{1}^{*},\dots,X_{m}^{*}$ with common distribution function equal to $\mathcal{M}\hat{F}_{n}^{d}$, consider a sample $\varepsilon_{1}^{*},\dots,\varepsilon_{m}^{*}$ from the same distribution as $\varepsilon$ and define $Y_{i}^{*}=X_{i}^{*}+\varepsilon_{i}^{*}$ for all $i$. Here, the sample $(\varepsilon_{i}^{*})$ is independent of all other variables $(Y_{1},\dots,Y_{n},X_{1}^{*},\dots,X_{m}^{*})$. The reason why we choose $m\ll n$ is explained in Section E.2.
@@
To study the properties of the test, we need to explore the limit behavior of the bootstrap version of the test statistic, both under the null and the alternative hypotheses. This is the aim of the following theorem, which is a bootstrap version of Theorem 5. A key ingredient in the proof of Theorem 5 is Theorem 2. Likewise, the proof of Theorem 6 relies on a bootstrap analogue of Theorem 2, namely Theorem 7, which is stated in Section C and proved in Section D. To prove that Assumptions B of that theorem holds in our setting, we rely on results from the broader deconvolution literature, that requires stronger assumptions than for Theorem 5. Specifically, we assume the following conditions D’ and E’ instead of D and E.
@@
where $\varphi_{m}^{*}(u)=m^{-1}\sum_{j=1}^{m}e^{iuY_{j}^{*}}.$ Moreover, we make similar assumptions as in Theorem 2 with $n$ replaced by $m$. However, in order to use results from [15] on convergence of estimators of moments in the deconvolution model (see Section E.4.2), we restrict attention to the specific kernel considered in that paper. Hence, we now use a kernel $K$ such that its Fourier transform is defined by
@@
where $r\geq 2$ is an even integer and $s\geq 1$ is an integer. In fact, we will assume that $r>4$ in order to ensure that $\hat{f}_{n}$ has a bounded absolute moment of order $2+\delta$, see Section E.4.4. Moreover, in order to use Corollary 1, we assume that $h_{n}^{2+2\beta}n\rightarrow\infty$ as $n\to\infty$. It follows from Lemma 6 below that the assumptions of Theorem 2 are satisfied under the assumptions of the following bootstrap version of it.
@@
    $T_{m}(F_{m}^{d*})\stackrel{{\scriptstyle L}}{{\longrightarrow}}\sup_{t\geq 0}|\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}(t)-\mathbb{G}_{F}(t)|$
@@
To understand the meaning of the conditional weak convergence in probability in (14), one can consider the Prohorov distance, that metrizes weak convergence of probability measures on separable spaces, see e.g. [4, page 72]. The Prohorov distance $\pi(\mu,\nu)$ between two probability measures $\mu$ and $\nu$ on $\ell^{\infty}(\mathbb{R}^{+})$ (the separable space that we consider in Theorem 6 above), is defined as the infimum of those $\epsilon>0$ for which the two inequalities
@@
$A^{\epsilon}=\left\{f\in\ell^{\infty}(\mathbb{R}^{+}),\ \inf_{g\in A}\|f-g\|_{\infty}<\epsilon\right\}.$
This distance metrizes weak convergence in the sense that a sequence of probability measures $(\mu_{n})_{n\in\mathbb{N}}$ weakly converges to a probability measure $\mu$ on $\ell^{\infty}(\mathbb{R}^{+})$ if and only if $\pi(\mu_{n},\mu)\to 0$ as $n\to\infty$. If we denote by $\mu_{n}^{*}$ the conditionnal distribution of $\sqrt{m}(F_{m}^{d*}-\hat{F}_{n}^{d})$ and by $\mu$ the distribution of $\mathbb{G_{F}}$, then the meaning of (14) is that
@@
Assume that $F$ is concave on $\mathbb{R}^{+}$. Then, it follows from Theorem 2 and Theorem 6 that the conditional limit distribution of $T_{m}(F_{m}^{d*})$ is the same as the limit distribution of $T_{n}(F_{n}^{d})$, so the test defined by the critical region in (12) satisfies the significance level constraint
$\lim_{n\to\infty}\mathbb{P}\left(T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right)=\gamma,$ (16)
provided that this limit distribution is continuous, see Section 23.1 in [32]. We will not attempt to rigourously prove that the limit distribution is continuous, insofar as this would constitute a separate research topic, but we give below some heuristics that are supported by our simulations below. In the case of weak concavity, where $F$ have at least an affine part, we conjecture that the limit distibution is continuous whence (16) holds. This is confirmed by numerical results in Section 6. On the other hand, if $F$ is strictly concave and nowhere affine, the limit of the test statistics is zero, by definition of $\mathcal{M}_{F}^{\prime}$. In particular, the limit distribution is not continuous. However, intuitively, the limit of the distance between an estimator and its least concave majorant tends to be stochastically larger if the underlying distribution function has affine parts (because, for an arbitrary concave function $F$, the value of the limiting distribution becomes larger if $F$ is replaced by an affine segment on a given interval where it is stricty concave). This is expected to occur more frequently in the boostrap case, where the underlying distribution function is $\mathcal{M}\hat{F}_{n}^{d}$, than in the original sample where the underlying distribution function $F$ is strictly concave. If these intuitions are correct, we may expect that the probability of rejection tends to 0, whence the test has asymptotic prescribed level $\gamma$, and is conservative in case of strict concavity. This is also confirmed by our simulation analysis in Section 6.

### v2

# arXiv:2601.04906v2[math.ST] 01 Mar 2026

@@ Abstract
We study nonparametric inference for a concave distribution function under the measurement error model, where the non-negative variable of interest is perturbed by additive independent noise. We propose a shape-constrained estimator defined as the least concave majorant on the non-negative real half-line of the deconvolution estimator of the cumulative distribution function, and we establish its uniform consistency as well as its square-root convergence in distribution. To assess the concavity assumption, we introduce a nonparametric test of the null hypothesis that the distribution function is concave on the non-negative real half-line against the alternative that it is not. The test is calibrated via the bootstrap. We show that the test statistic and its bootstrap analogue have the same limiting distribution under the null, while the rejection probability tends to one under the alternative. The proofs mainly rely on a bootstrap Donsker-type result for the deconvolution estimator of the cumulative distribution function, combined with the functional delta method. Simulation studies illustrate the finite-sample performance of both the estimator and the test.
@@ 1 Introduction
Statistical inference under shape restrictions exploits nonparametric assumptions on the shape of the distribution of interest to improve estimates. Additionally, it deals with nonparametric tests about such shape constraints. The books by [35, 23, 40] provide an overview of the issues addressed in the context of statistical inference under shape constraints. The results depend on the methodology and on the type of assumption, see for instance [22, 31, 15, 36, 26, 28, 27]. One of the most commonly used constraints, starting from the seminal work of [22], is the concavity of the cumulative distribution function [9, 5], corresponding to a decreasing density [23, Chapter 2.2].
@@
This paper combines theories on deconvolution estimation (i.e., estimation under measurement error) and on estimation and testing under shape constraints. As a shape-constrained estimator, we consider the least concave majorant of the classic kernel deconvolution estimator of the distribution function $F$ of $X$. We show its uniform consistency under general assumptions, and we establish a Donsker-type result for it, i.e. we show that the difference between the estimator and the true concave distribution function converges weakly at the rate $\sqrt{n}$ to a centered Borel random variable in $\ell_{\infty}(\mathbb{R})$. This latter result exploits the results of [41] combined with the functional delta method.
Then, in order to test concavity of $F$, we consider the test statistic defined as the supremum distance between the kernel deconvolution estimator and its least concave majorant. A similar approach has been developed in several settings to test concavity of a given function (or monotonicity of its derivative when it exists). Depending on the setting, the function of interest could be, for instance, a cumulative hazard rate, a distribution function, a copula function. The general idea is to define the test statistic as the $L_{p}$-distance, with some fixed $p\in[1,\infty]$, between an unconstrained estimator of the target function and its least concave majorant. Two different strategies have been proposed to calibrate such a test. The first one relies on a least favorable hypothesis, as discussed in [17, 15, 25, 8, 13, 3]. This approach consists in showing that a linear target function is least favorable in the sense that the test statistic is (at least asymptotically) stochastically larger under linearity than it would be under a concave but nonlinear function. The limiting distribution is then derived under this least favorable case, and the corresponding $(1-\gamma)$-quantile is used as the critical value, ensuring that the test achieves the desired asymptotic significance level $\gamma$. The second strategy consists in deriving the limit distribution of the test statistic under every arbitrary concave functions and to approximate the limit using either plug-in or bootstrap. This gives more powerful tests than those based on a least favorable hypothesis. However, a major difficulty with this approach stems from the fact that the rate of convergence of the test statistic differs depending on whether or not the target function is strictly concave. For instance in case of a distribution function on $[0,1]$, it is of order $n^{1/2}$ under linearity and $n^{2/3}$ under strict concavity with appropriate smoothness conditions, see [25]. The latter paper points out that, due to different rates of convergence, scaling the test statistic by $n^{2/3}$ necessarily leads to a test that does not have the prescribed asymptotic level. On the other hand, scaling the test statistic by $n^{1/2}$ leads to a degenerate limit distribution under strict concavity with no affine segments. This makes it difficult to prove the consistency of a bootstrap calibration since such proofs rely on continuity of the limit distribution, see e.g. [43, Chapter 23]. Moreover, when non degenerate, the limiting distribution is too complex to be feasibly approximated using plug-in methods. In three different settings, [37, 4, 20] develop such a methodology, and prove that their test statistic and its bootstrap version share the same limit distribution under the null hypothesis. Hence, the asymptotic probability of rejection is the prescribed level under each concave function for which the limit distribution of the corresponding test statistic is continuous. It is unclear what is the asymptotic probability under strict concavity with no affine parts since the limit distribution degenerates at zero in that case. [37, 4] have neither simulations nor theoretical results for that case. On the other hand, [20] slightly modifies the test statistic to cover some cases of strict concavity albeit some other cases are not covered, and simulations show good performance even under some cases of strict concavity.
In the model with measurement error, we do not have access to a least favorable hypothesis (see Section 5.3) so we scale the test statistic by $n^{1/2}$ and consider a bootstrap calibration. In the same spirit as in [37, 4, 20], computing the limit distribution of the rescaled test statistic under arbitrary concave function relies on computing the limit distribution of the unconstrained estimator (provided in our setting by [41]), and combine it with Hadamard directional differentiability of the least concave majorant operator and with the delta method. We build a boostrap statistic that differs in spirit from those in [37, 4, 20], and that has the same limit distribution as the original one under any concave distribution function. The bootstraps used in the three papers we have just cited use a similar strategy, inspired by [14, 19], to circumvent the fact that the standard bootstrap does not work, due to the very weak form of differentiability of the least concave majorant operator. For our part, we circumvent this difficulty by “sub-sampling” in the sense that our bootstrap sample has size $m\ll n$. Moreover, whereas the above papers sample from a sort of empirical distribution and then adapt their bootstrap test statistic so that it mimics the behaviour the original one would have under the null hypothesis, we directly sample from a distribution that satisfies the null hypothesis. We show that if the distribution function is concave but not strictly concave (meaning that it has at least an affine part), the probability of rejection converges to the prescribed level as the sample size goes to infinity. On the other hand, the common limit distribution is again degenerate at zero under strict concavity with no affine segments. We provide some heuristics on the fact that the asymptotic probability of rejection in that case is smaller than the prescribed level, and we show simulation experiments in which the probability becomes very close to zero as the sample size increases.
The performance of the constrained estimator and that of the test are illustrated through a simulation study. The simulations suggest that, for sufficiently large sample sizes and a moderate noise-to-signal ratio, (i) when $F$ is concave, the constrained estimator outperforms the unconstrained one; (ii) under concavity of $F$, the rejection rate of the test does not exceed the precribed level; and (iii) the test can detect nonconcavity of $F$ even in settings where the distribution of $Y$ is concave.
@@
    a bootstrap version of the Donsker Theorem 1 in [41] from which, as a particular case, one can derive the limit behavior of the deconvolution estimator of the distribution function in a bootstrap setting, under measurement error, see Theorem 8 in Section A and (14) below;
@@
The paper is organized as follows. Section 2 provides the model, definitions and assumptions. The constrained estimator is studied in Section 3, which also contains new results about uniform convergence of least concave majorants of continuously differentiable functions, as well as that of their slopes. The test statistic is defined in Section 4, where its limit behavior is studied under both the null and the alternative hypotheses. A bootstrap calibration of the test is given in Section 5, which also contains an alternative calibration that provides a less powerfull test, compared to the bootstrap, but is easier to implement. The main part of the paper ends with Section 6, which is devoted to simulations. The main sections are followed by appendices that are organized as follows. Section A provides a generalization of the Donsker result in [41] to a bootstrap setting. The results from Sections 2 and 3 are proved in Section B. Theorem 5, that provides the limit behavior of the test statistic, is proved in C, while the proofs of Theorem 6 and Theorem 7, that prove the validity of bootstrap calibration, are given in Section D. The results from Section A are proved in Section E.
@@ 2.2 Model, objective and basic estimators
Our aim is to obtain a test for the null hypothesis
@@
against the alternative hypothesis $\mathcal{H}_{1}:\ ``F\text{ is not concave on }[0,\infty)$”. Moreover, if the null hypothesis is not rejected, we wish to construct an estimator of $F$ that takes it into account, i.e. a concave estimator. Defining the concave estimator and the test requires to first define Fourier transforms, which we will now do.
@@ 2.3 Assumptions
A key ingredient in our proofs is Theorem 1 in [41]. This is a Donsker-type result that gives, as a special case, the limiting behavior of the deconvolution estimator of the distribution function in the measurement error model. Assumptions K, D and E below are taken from that theorem.
@@ 3.1 Definition of the concave estimator and consistency
The following theorem, which is proved in Section B, provides a general result on convergence of least concave majorants (and their slopes) of a sequence of continuously differentiable functions whose sequence of slopes uniformly converge.
@@
As a corollary (also proved in Section B) of Theorem 1, we derive uniform convergence of our concave estimators.
@@ 3.2 Limit distribution
Theorem 2 ([41]).
Assume conditions D and E, with $\beta<1/2$, $\alpha+3\gamma>2\beta+1$ for some $\gamma\in(\beta,1/2)$ and K with $L=\left\lfloor\alpha+\gamma\right\rfloor$. Furthermore, let $h_{n}^{2\alpha+2\gamma}n\rightarrow 0$ and $h_{n}^{\rho}n\rightarrow\infty$ for some $\rho>4\beta-4\gamma+2$. Then, we have
@@
We are now in position to state the main result of the section, which gives the limit distribution of the concave estimator $\mathcal{M}F_{n}^{d}$. The theorem shows that the estimator converges to the least concave majorant of $F$ at the rate $\sqrt{n}$, so, in particular, the estimator converges to $F$ if the latter is concave on its support. More precisely, the difference between the estimator and the least concave majorant of $F$ converges at the rate $\sqrt{n}$ to a centered Borel variable in $\ell_{\infty}(\mathbb{R})$.
@@
$\sqrt{n}(\mathcal{M}F_{n}^{d}-\mathcal{M}F)\rightsquigarrow\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}(0)\quad in\ \ell_{\infty}(\mathbb{R}^{+}).$ (9)
@@ 5.1 Bootstrap calibration
We turn to the construction of the bootstrap observations. Fix $m\ll n$ that tends to infinity as $n\to\infty$, and conditionally on the original sample $Y_{1},\dots,Y_{n}$, consider a sample of $m$ iid random variables $X_{1}^{*},\dots,X_{m}^{*}$ with common distribution function equal to $\mathcal{M}\hat{F}_{n}^{d}$, consider a sample $\varepsilon_{1}^{*},\dots,\varepsilon_{m}^{*}$ from the same distribution as $\varepsilon$ and define $Y_{i}^{*}=X_{i}^{*}+\varepsilon_{i}^{*}$ for all $i$. Here, the sample $(\varepsilon_{i}^{*})$ is independent of all other variables $(Y_{1},\dots,Y_{n},X_{1}^{*},\dots,X_{m}^{*})$. The reason why we choose $m\ll n$ is explained in Section D.
@@
To study the properties of the test, we need to explore the limit behavior of the bootstrap version of the test statistic, both under the null and the alternative hypotheses. This is the aim of the following theorem, which is a bootstrap version of Theorem 5. A key ingredient in the proof of Theorem 5 is Theorem 2. Likewise, the proof of Theorem 6 relies on a bootstrap analogue of Theorem 2, namely Theorem 8, which is stated in Section A and proved in Section E. To prove that Assumptions B of that theorem holds in our setting, we rely on results from the broader deconvolution literature, that requires stronger assumptions than for Theorem 5. Specifically, we assume the following conditions D’ and E’ instead of D and E.
@@
where $\varphi_{m}^{*}(u)=m^{-1}\sum_{j=1}^{m}e^{iuY_{j}^{*}}.$ Moreover, we make similar assumptions as in Theorem 2 with $n$ replaced by $m$. However, in order to use results from [24] on convergence of estimators of moments in the deconvolution model (see Section D.5.2), we restrict attention to the specific kernel considered in that paper. Hence, we now use a kernel $K$ such that its Fourier transform is defined by
@@
where $r\geq 2$ is an even integer and $s\geq 1$ is an integer. In fact, we will assume that $r>4$ in order to ensure that $\hat{f}_{n}$ has a bounded absolute moment of order $2+\delta$, see Section D.5.4. Moreover, in order to use Corollary 1, we assume that $h_{n}^{2+2\beta}n\rightarrow\infty$ as $n\to\infty$. It follows from Lemma 3 below that the assumptions of Theorem 2 are satisfied under the assumptions of the following bootstrap version of it.
@@
    $T_{m}(F_{m}^{d*})\stackrel{{\scriptstyle L}}{{\longrightarrow}}\|\mathcal{M}_{F}^{\prime}\mathbb{G}_{F}-\mathbb{G}_{F}\|_{\infty}$
@@
To understand the meaning of the conditional weak convergence in probability in (14), one can consider the Prohorov distance, that metrizes weak convergence of probability measures on separable spaces, see e.g. [6, page 72]. The Prohorov distance $\pi(\mu,\nu)$ between two probability measures $\mu$ and $\nu$ on $\ell_{\infty}(\mathbb{R}^{+})$ (the separable space that we consider in Theorem 6 above), is defined as the infimum of those $\epsilon>0$ for which the two inequalities
@@
$A^{\epsilon}=\left\{f\in\ell_{\infty}(\mathbb{R}^{+}),\ \inf_{g\in A}\|f-g\|_{\infty}<\epsilon\right\}.$
This distance metrizes weak convergence in the sense that a sequence of probability measures $(\mu_{n})_{n\in\mathbb{N}}$ weakly converges to a probability measure $\mu$ on $\ell_{\infty}(\mathbb{R}^{+})$ if and only if $\pi(\mu_{n},\mu)\to 0$ as $n\to\infty$. If we denote by $\mu_{n}^{*}$ the conditionnal distribution of $\sqrt{m}(F_{m}^{d*}-\hat{F}_{n}^{d})$ and by $\mu$ the distribution of $\mathbb{G_{F}}$, then the meaning of (14) is that
@@
5.2 Asymptotic properties of the test
We assume without loss of generality that $\gamma<1/2$ since levels of tests are always choosen with this constraint (recall that a typical level is $\gamma=0.05$). The following theorem, which is proved in Section D.3, shows that the test has the prescribed asymptotic level $\gamma$ at each point $F$ in the null hypothesis that has at least an affine part. The proof of the theorem uses the fact that, with probability one, the process $\mathbb{G}_{F}$ cannot be affine on a given non-degenerate interval on which $F$ is affine. We show in Section D.3 that this property holds in the case where the distribution of $\epsilon$ is supported on the whole real line, hence we make this assumption in the theorem. We conjecture that this holds under more general conditions, but due to the very complex form of the covariance function of $\mathbb{G}_{F}$, proving it is out of the scope of the paper. We can notice that in the case of zero noise, the limit process is $B\circ F$ where $B$ is standard Brownian bridge, and it is well known that with probability one, $B$ (hence also $B\circ F$) cannot be affine on a given interval on which $F$ is affine. Hence, the conjecture generalizes to the case of measurement error a property that is well know under no error.
Theorem 7.
Assume $\gamma<1/2$, the distribution of $\epsilon$ is supported on $\mathbb{R}$, and the assumptions of Theorem 6 are satisfied. If $F$ is concave on $\mathbb{R}^{+}$ with at least an affine part, then
$\lim_{n\to\infty}\mathbb{P}\left(T_{n}(F_{n}^{d})>c_{\gamma,m}^{*}\right)=\gamma.$ (16)
On the other hand, if $F$ is strictly concave and nowhere affine, the limit of the test statistic is zero. In particular, the limit distribution is not continuous and we doubt that (16) holds. In fact, our conjecture is that the limit is equal to zero. The conjecture is based on the following heuristics. Since the original test statistic $T_{n}(F_{n}^{d})$ converges to zero in probability, it is expected that there exists a sequence of positive numbers that tend to infinity and such that $v_{n}T_{n}(F_{n}^{d})$ is stochastically bounded and bounded away from zero with hight probability. In such a case, one can expect that the bootstrap test statistic has a similar behaviour whence $v_{m}T_{m}(F_{n}^{d*})$ is bounded from above and away from zero with hight probability. Then, its $(1-\gamma)$-quantile is of order of magnitude $v_{m}^{-1}$. Since $m\ll n$ one has $v_{m}^{-1}\gg v_{n}^{-1}$ whence the probability that $T_{n}(F_{n}^{d})$, which is of order $v_{n}^{-1}$, exceeds the quantile of order $v_{m}^{-1}$ tends to zero as $n$ tends to infinity. If these intuitions are correct, we may expect that the probability of rejection tends to 0, whence the test has asymptotic prescribed level $\gamma$ at any point $F$ in the null hypothesis, and is conservative in case of strict concavity. This is confirmed by our simulation analysis in Section 6.